\begin{aligned}π & = arctan (\frac{b}{d}) \\ π & ≃ (\frac{b}{d}) \end{aligned}

d = (\frac{b}{π} )

π = 100 ± 20 mas

π = 10 ± 7 mas

fractional uncertainty in parallax = \frac{σ_{π}}{π}

\frac{10 mas}{π} > 0.10 → π > 100 mas
→ d < 10 pc

π = 2.39 ± 0.03 mas

size of PSF θ ∼ \frac{λ}{D}